McCoy Brothers manufactures and sells two products, A and Z in the ratio of 4:2. Product A sells for $75; Z sells for $95. Variable costs for product A are $35; for Z $40. Fixed costs are $418,500. Compute the number of units of Product Z McCoy must sell to break even.

Multiple Choice

  • 9,300.

  • 6,200.

  • 1,550.

  • 3,100.

  • 6,750.
Answer #1

Contribution Margin

Of A = 75-35 = 40

Of Z = 95-40 = 55

Weighted average contribution margin

= (40*4/6) + (55*2/6)

= 45

Breakeven sales = Fixed cost /Weighted average contribution margin

= 418,500/45

= 9300